OpenStudy (anonymous):

23. Which is corect regarding the statement: “If x is an odd integer, then the median of x, x + 2, x + 6, and x + 10 is an odd number” there is not enough information provided to answer the question the statement is always true the statement is sometimes true the statement is always false

OpenStudy (anonymous):

So the median of this data set is the mean of the two middle numbers, which are x+2 and x+6 Mean = ((x+2)+(x+6))/2 Mean = (2x+8)/2 Mean = x+4 We can say from this addition that if x is an odd number then the mean will also always be odd. By extension we can therefore say that the median of the set will also be odd. Hence the answer is B

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

You're welcome. If you like my help then feel free to give a medal using the best response button to the right of this message :)